Varmit-B-Gone is a pest control service that operates in a suburban neighborhood. The company attempts to make service calls at least once a month to all homes that subscribe to its service. It makes more frequent calls during the summer. The number of subscribers also varies with the season. The number of subscribers and the average number of calls to each subscriber for the months of interest follow:

Subscribers Service calls (per subscriber)
March 400 0.6
April 600 0.9
May 1,450 1.4
June 1,650 2.3
July 1,600 2.7
August 1,550 2.8


The average price charged for a service call is $40. Of the service calls, 30 percent are paid in the month the service is rendered, 60 percent in the month after the service is rendered, and 8 percent in the second month after. The remaining 2 percent is uncollectible.


Required:

What are Varmit-B-Gone’s expected cash receipts for May, June, July, and August?
Answer #1

First, we can calculate service call revenues for all 6 months

March April May June July August
Number of subscribers (a) 400 600 1450 1650 1600 1550
Service call per subscriber (b) 0.6 0.9 1.4 2.3 2.7 2.8
Total service calles (c = a*b) 240 540 2030 3795 4320 4340
Average price per service call (d) $ 40 $ 40 $ 40 $ 40 $ 40 $ 40
Total Service call revenues (e = c*d) $ 9,600 $ 21,600 $ 81,200 $ 151,800 $ 172,800 $ 173,600

Now we can go for expected cash collections

May June July August
Cash received from March service calls $9,600*8% = $768
Cash received from April service calls $21,600*60% = $12,960 $21,600*8% = $1,728
Cash received from May service calls $81,200*30% = $24,360 $81,200*60% = $48,720 $81,200*8% = $6,496
Cash received from June service calls $151,800*30% = $45,540 $151,800*60% = $91,080 $151,800*8% = $12,144
Cash received from July service calls $172,800*30% = $51,840 $172,800*60% = $103,680
Cash received from August service calls $173,600*30% = $52,080
Expected cash collections $                                 38,088 $                                   95,988 $                                 149,416 $                                   167,904

You can reach me over comment box if you have any doubts. Please rate this answer